the length of a rectangular playing field is 5ft less than twice its width. If the perimeter of the playing field is 230ft, find the length and width on the field.
Answer by ptaylor(2198)

Let x=width (w)
Then 2x-5=length(l)
Perimeter=230=2w+2l So, our equation to solve:
2x+2(2x-5)=230 get rid of parens
2x+4x-10=230 add 10 to both sides
2x+4x-10+10=230+10 collect like terms
6x=240 divide both sides by 6
x=40 ft--------------width
2x-5=80-5=75 ft ------------length
ck
2(75)+2(40)=230
150+80=230
230=230

Hope this helps----ptaylor